# ZIP-FIT: Embedding-Free Data Selection via Compression-Based Alignment

### v1

@@ 2.2 ZIP-FIT Algorithm
10:Select the top $k$ examples from $D$ based on the highest alignment scores.
@@ 4 Higher Alignment leads to more efficient training
Figure 4: Highly aligned data lowers cross-entropy loss more efficiently. The x-axis shows the number of training tokens, and the y-axis represents the cross-entropy (CE) test loss. Different curves correspond to datasets filtered by different alignment scores, indicating their relevance to the target domain. The most aligned data reduce Test CE loss significantly faster than less aligned data. The left panel depicts results using GPT-2, and the right panel uses Mistral7B, demonstrating that using highly aligned data not only accelerates training but also achieves better model performance, validating the effectiveness of ZIP-FIT for data selection in fine-tuning.
@@ 6 Impact of Data Misalignment on Model Performance
Experiment: We fine-tuned the Mistral7B model on the same source dataset we used for the AutoFormalization experiment (see Appendix 5.1), filtering data with ZIP-FIT at different alignment thresholds ($>$0.1, $>$0.2, $>$0.3). Each threshold creates a progressively more aligned dataset, where the $>$0.3 dataset is the most aligned, and the $>$0.2 dataset is a superset of the $>$0.3 dataset, including less aligned data. Similarly, the $>$0.1 dataset is a superset of both $>$0.2 and $>$0.3. Figure 6 shows cross-entropy test loss (y-axis) versus the number of training tokens (x-axis). The dashed line marks the pretrained Mistral7B baseline.
Results: ZIP-FIT selected data achieves lower cross-entropy loss faster than training on all data, showing improved performance and efficiency. Higher alignment thresholds result in a steeper loss reduction, confirming that filtering out misaligned data enhances fine-tuning.Misalignment in training data can introduce noise and irrelevant patterns, which typically require more training data and computational resources to overcome. By applying higher alignment thresholds, ZIP-FIT ensures that only the most relevant and helpful examples are used for training. This targeted selection leads to a more efficient learning process as evidenced by the sharper decline in cross-entropy loss for higher alignment thresholds. Such efficiency is crucial in scenarios where computational resources are limited or costly.
@@
Future Directions: Further research could explore adaptive alignment thresholds based on real-time validation performance, potentially automating the selection process to optimize both speed and accuracy during training.
These results further validate the empirical performance gains and computational efficiency achieved by ZIP-FIT, as outlined in our contributions. By filtering out misaligned data, ZIP-FIT accelerates fine-tuning and reduces computational overhead, confirming its utility in low-resource settings.
@@ 7 Related Works
Curating pre-training data for Language Models often involves using classifiers to filter high-quality data from large corpora like Common Crawl, as done for models like GPT-3 and PaLM2 (Brown et al., 2020; Google, 2023; Shao et al., 2024). While effective, this process requires significant computational resources and large volumes of curated data. In contrast, ZIP-FIT efficiently selects relevant data without relying on external models, making it especially useful in data-scarce environments.
@@
Mixture weights are essential when drawing from multiple domains, as they significantly influence the performance of language models (Du et al., 2022; Xie et al., 2023b). While Domain Reweighting with Minimax Optimization DoReMi) (Xie et al., 2023a) proposes a robust domain-level reweighting strategy suitable for diverse target distributions. DoReMi is not designed for example-level data selection, as it primarily focuses on domain-level reweighting. Adapting it to select individual data points for specific target distributions would require substantial modifications to its foundational algorithm. One possibility be to effectively transform each data point into a ’mini-domain,’ a process that would stray significantly from DoReMi’s original purpose and scope. Therefore, we did not use DoReMi in our comparisons because it does not directly address the fine-grained selection needs that ZIP-FIT fulfills.
@@ 9 Discussion and Future Work
ZIP-FIT introduces an efficient, embedding-free approach for data selection in language model fine-tuning. By leveraging gzip compression to capture redundancies in data, ZIP-FIT enables the alignment of large-scale datasets with a target domain without the computational burden of neural embeddings. Our results show that using compression-based alignment leads to faster convergence and lower cross-entropy loss compared to existing methods like DSIR and D4 (Tirumala et al., 2023; Xie et al., 2023b).
@@ Appendix C Additional Experimental Results: data selection for efficient fine-tuning using ZIP-FIT
Appendix D Data Selection Profiling (Run Times)
@@
Appendix E Qualitative Analysis
@@
Appendix F Future Work (Cont.)

### v2

@@ 2.2 ZIP-FIT Algorithm
10:Select the top-K examples from $D$ based on the highest alignment scores.
@@ 4 Higher Alignment leads to more efficient training
Figure 4: Highly aligned data lowers cross-entropy loss more efficiently. The x-axis shows the number of training tokens, and the y-axis represents the cross-entropy (CE) test loss on the ProofNet test set. Different curves correspond to datasets filtered by different alignment scores, indicating their relevance to the target domain. The most aligned data reduce Test CE loss significantly faster than less aligned data. The left panel depicts results using GPT-2, and the right panel uses Mistral7B, demonstrating that using highly aligned data not only accelerates training but also achieves better model performance, validating the effectiveness of ZIP-FIT for data selection in fine-tuning.
@@ 6 Impact of Data Misalignment on Model Performance
Experiment: We fine-tuned the Mistral7B model on the same source dataset we used for the AutoFormalization experiment (see Appendix 5.1), filtering data with ZIP-FIT at different alignment thresholds ($>$0.1, $>$0.2, $>$0.3). Each threshold creates a progressively more aligned dataset, where the $>$0.3 dataset is the most aligned, and the $>$0.2 dataset is a superset of the $>$0.3 dataset, including less aligned data. Similarly, the $>$0.1 dataset is a superset of both $>$0.2 and $>$0.3. Figure 6 shows CE test loss (y-axis) versus the number of training tokens (x-axis).
Results: ZIP-FIT selected data achieves lower CE loss faster than training on all data (Figure 6), showing improved performance and efficiency. Higher alignment thresholds result in a steeper loss reduction, confirming that filtering out misaligned data enhances fine-tuning. Misalignment can introduce noise and irrelevant patterns, which we hypothesize require more training data and computational resources to overcome. Applying higher alignment thresholds, ZIP-FIT ensures that only the most relevant examples are used for training. This targeted selection leads to a more efficient learning process as evidenced by the sharper decline in CE loss for higher alignment thresholds. Such efficiency is crucial in scenarios where computational resources are limited or costly.
@@
Future Directions: Could explore adaptive alignment thresholds based on real-time validation CE, potentially automating the selection process to optimize both speed and accuracy during training.
Filtering out misaligned data accelerates fine-tuning and reduces computational overhead, confirming its performance gains and computational efficiency as outlined in our contributions.
@@ 7 Related Works
Curating pre-training data often involves using classifiers to filter high-quality data from large corpora like Common Crawl, as done for models like GPT-3 and PaLM2 (Brown et al., 2020; Google, 2023; Shao et al., 2024). While effective, this process requires significant computational resources and large volumes of curated data. In contrast, ZIP-FIT efficiently selects relevant data without relying on external models, making it especially useful in data-scarce environments.
@@
Mixture weights are essential when drawing from multiple domains, as they influence the performance of language models (Du et al., 2022; Xie et al., 2023b). DoReMi (Domain Reweighting with Minimax Optimization) (Xie et al., 2023a) proposes a reweighting strategy suitable for handling diverse target distributions, but it primarily focuses on adjusting weights at the domain level. Adapting it to select individual data points for specific target distributions would require substantial modifications to its foundational algorithm. One potential approach would be to effectively transform each data point into a ’mini-domain,’ a process that would stray significantly from DoReMi’s original purpose and scope. Therefore, we did not use DoReMi in our comparisons because it does not directly address the fine-grained selection needs that ZIP-FIT fulfills.
@@
Compression: (Pandey, 2024) demonstrates that gzip-compressibility predicts how language model scaling laws shift with data complexity, challenging the data-agnostic assumptions of Chinchilla-style scaling laws. (Jiang et al., 2022) use gzip-based compression distance with kNN for zero-shot text classification, outperforming BERT on some datasets. (Delétang et al., 2024) show that language modeling is equivalent to compression, where even gzip can define predictive distributions via coding length. (Yoran et al., 2025) propose the KoLMogorov Test, showing that gzip remains a strong baseline for code-based compression, outperforming pretrained LLMs on real data. show that gzip compression ratio is a fast, effective proxy for text diversity, capturing key repetition patterns in LLM outputs and aligning with slower lexical metrics e.g., self-BLEU.
@@ 9 Discussion and Future Work
ZIP-FIT introduces an efficient, embedding-free approach for data selection in language model fine-tuning. By leveraging compression algorithms to capture redundancies in data, ZIP-FIT enables the alignment of large-scale datasets with a target domain without the computational burden of neural embeddings. Our experiments with different compression algorithms (Figure 8) reveal that lighter compression (e.g., LZ4 at level 0) leads to better performance, achieving a 12.19% Pass@1 on HumanEval compared to 11.58% with gzip. This suggests that while compression effectively captures alignment signals, aggressive compression can remove subtle but important patterns. These insights highlight the importance of careful selection of compression parameters in optimizing the quality of data selection. Our results show that using compression-based alignment leads to faster convergence and lower cross-entropy loss compared to existing methods like DSIR and D4 (Tirumala et al., 2023; Xie et al., 2023b).
@@ Appendix C Additional Experimental Results: data selection for efficient fine-tuning using ZIP-FIT
Appendix D Baseline Comparison using Pass@1 on HumanEval
ZIP-FIT demonstrates substantial improvements in code generation capability across different fine-tuning approaches. When applied to 4-bit quantized models with LoRA fine-tuning, ZIP-FIT doubles the Pass@1 performance of the base model and outperforms existing data selection methods.
Table 1: Performance and efficiency comparison of data selection methods. Results show Pass@1 and Pass@10 scores on HumanEval using top 1M tokens for fine-tuning, along with data selection time. Data selection times exclude fine-tuning time.
Fine-tuning Data Selection Pass@1 (%) Pass@10 (%) Selection Time
None Pre-trained Gemma2-2B 15.24 38.81
None Pre-trained Gemma2-2B (4-bit quantized) 6.09
Full FT ZIP-FIT 18.86 41.78 32s
Full FT LESS 18.06 40.19 19h
Full FT DSIR 17.98 44.27 97s
QLoRA ZIP-FIT 12.19 32s
QLoRA DSIR 9.14 97s
QLoRA D4 6.09 7h 40m
Appendix E Impact of Compression Algorithms and Levels
Figure 8: Lighter compression preserves crucial information for data selection. At minimum compression levels, both gzip and LZ4 achieve the strongest Pass@1 scores (11.58% and 12.19%), significantly outperforming the base model (6.09%, dashed line). Performance systematically degrades with increased compression across all algorithms, suggesting that aggressive compression removes valuable alignment signals.
To investigate the impact of different compression algorithms on ZIP-FIT’s performance, we conducted experiments comparing three widely used compression methods: gzip, zstd, and LZ4. Each algorithm was tested across its available compression levels, normalized to a 0-1 scale for comparison. As shown in Figure 8, compression algorithm choice and level significantly impact performance.
Key findings include:
  • LZ4 at minimum compression achieves the best performance (12.19% Pass@1)
  • Higher compression levels generally lead to decreased performance across all algorithms
  • gzip shows more stable performance degradation compared to LZ4 and zstd
  • zstd consistently underperforms relative to both GZIP and LZ4
These results suggest that lighter compression better preserves the structural information needed for effective data selection. The superior performance of LZ4 at minimal compression indicates that aggressive data compression may remove subtle but important patterns useful for alignment assessment.
Appendix F Data Selection Profiling (Run Times)
@@
Appendix G Qualitative Analysis
@@
Appendix H Future Work (Cont.)
@@
Autonomous Validation Set Generation: A current limitation of ZIP-FIT is its dependence on a small, curated validation set (e.g., 185 samples for ProofNet and 82 samples for half the HumanEval test set). Future work could explore the use of generative models to create synthetic validation sets from task-specific instructions. This approach could also be expanded to enable autonomous self-directed, model-driven generation of validation data.